OpenStudy (anonymous):

Simplify (4 + √3)^2 Answers: 19 + 8√3 16+2√3 19

OpenStudy (anonymous):

its just like how you do (x+ y)^2

OpenStudy (anonymous):

let 4 be the x and\[\sqrt{3}\] be y :)

OpenStudy (owlcoffee):

I'll help ya: we know that to expand a squared bin. we have this little formula: \[(a+b)^{2}=a ^{2}+2ab+b ^{2}\] Let's apply it to the problem: \[(4+\sqrt{3})^{2}=(4)^{2}+2(4)(\sqrt{2})+(\sqrt{3})^{2}\] Okay, let's operate: \[(16)+2(4)(\sqrt{3})+3\] All i did was squaring the 4, and since Sqrt and exponential are opposite ican simplify them. Now, we know that 4= Sqrt of 16 and 2 is sqrt of 4, let's apply it: \[16+\sqrt{4}(\sqrt{16})(\sqrt{3})+3\] Knowing a little radication, we get: \[16+\sqrt{(4)(16)(3)}+3\] we can simplify it as: \[19+8\sqrt{3}\] And done.

OpenStudy (anonymous):

(4+sqrt(3))^2=16+3+8sqrt(3)=19+8 sqrt(3)